Two planes started at the same time from the same airport and flew in opposite directions. One flew 60 miles per hour faster than the other. In 5 hours, they were 2800 miles apart. Find the rate of each plane.
Found 2 solutions by whty89, Paul:
Answer by whty89(20)

This question is too hard

Answer by Paul(988)

Sicne one flew 60 miles faster than the other we write 60+x
And for the slower we write x
In 5 hours they flew 2800 so 5 times the sum of two planes

5(60+x+x)=2800
300+10x=2800
10x=2500
x=250

250+60=310

Hence one plane flew at 250mph and the other flew at 310mph.
Paul.